### v1

# MathFusion: Enhancing Mathematic Problem-solving of LLM through Instruction Fusion

Abstract
Large Language Models (LLMs) have shown impressive progress in mathematical reasoning. While data augmentation is promising to enhance mathematical problem-solving ability, current approaches are predominantly limited to instance-level modifications—such as rephrasing or generating syntactic variations—which fail to capture and leverage the intrinsic relational structures inherent in mathematical knowledge. Inspired by human learning processes, where mathematical proficiency develops through systematic exposure to interconnected concepts, we introduce MathFusion, a novel framework that enhances mathematical reasoning through cross-problem instruction synthesis. MathFusion implements this through three fusion strategies: (1) sequential fusion, which chains related problems to model solution dependencies; (2) parallel fusion, which combines analogous problems to reinforce conceptual understanding; and (3) conditional fusion, which creates context-aware selective problems to enhance reasoning flexibility. By applying these strategies, we generate a new dataset, MathFusionQA, followed by fine-tuning models (DeepSeekMath-7B, Mistral-7B, Llama3-8B) on it. Experimental results demonstrate that MathFusion achieves substantial improvements in mathematical reasoning while maintaining high data efficiency, boosting performance by 18.0 points in accuracy across diverse benchmarks while requiring only 45K additional synthetic instructions, representing a substantial improvement over traditional single-instruction approaches. Our datasets, models, and code are publicly available at https://github.com/QizhiPei/mathfusion.
MathFusion: Enhancing Mathematic Problem-solving of LLM
through Instruction Fusion
Qizhi Pei1, Lijun Wu2∗, Zhuoshi Pan3, Yu Li2, Honglin Lin2, Chenlin Ming4, Xin Gao2, Conghui He2∗, Rui Yan1,5††thanks: Corresponding authors: Lijun Wu (wulijun@pjlab.org.cn), Conghui He (heconghui@pjlab.org.cn), and Rui Yan (ruiyan@ruc.edu.cn). 1Gaoling School of Artificial Intelligence, Renmin University of China 2Shanghai AI Laboratory 3Tsinghua University 4Shanghai Jiao Tong University 5School of Computer Science, Wuhan University {qizhipei,ruiyan}@ruc.edu.cn {wulijun,heconghui}@pjlab.org.cn
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in various reasoning tasks (Wei et al., 2022; Huang and Chang, 2023), with mathematical problem-solving emerging as a critical domain for assessing their cognitive abilities (Ahn et al., 2024).
Specialized mathematical LLMs have emerged to address the unique challenges of solving complex mathematical problems (Yang et al., 2024; Shao et al., 2024; Ying et al., 2024; team, 2024). Current approaches to enhance mathematical reasoning primarily focus on four paradigms: continued pre-training with math corpora (Yang et al., 2024; Shao et al., 2024), reinforcement learning (RL) from human or automated feedback (Luo et al., 2023; Lu et al., 2024), test-time compute scaling (Wang et al., 2024a; Kang et al., 2024; Guan et al., 2025; Xi et al., 2024), and supervised fine-tuning (SFT) using problem-solution pairs (Tang et al., 2024; Tong et al., 2024). Among these, SFT is the most widely adopted paradigm (Setlur et al., 2024) due to its simplicity. However, its effectiveness is often limited by the complexity and diversity of the mathematical training data (Luo et al., 2023) during SFT. To this end, data augmentation and synthesis have emerged as promising directions to enhance mathematical reasoning. For example, approaches such as MetaMath (Yu et al., 2024) and WizardMath (Luo et al., 2023) emphasize enhancing individual problems through rephrasing and difficulty variation.
While instance-level modifications have shown potential, they do not resolve the fundamental challenge: the inability of LLMs to effectively capture and leverage the intrinsic relational structures that characterize mathematical knowledge (Chu-Carroll et al., 2024; Srivatsa and Kochmar, 2024). This limitation becomes particularly apparent in real-world scenarios, where complex mathematical problems are often composed of interdependent sub-problems that form intricate dependency graphs (Bagherzadeh et al., 2019; Prabawa et al., 2023). For instance, solving a system of equations requires the sequential solution of individual equations, followed by the reconciliation of constraints.
Motivated by the way human learners develop proficiency through systematic exposure to interconnected ideas (Komarudin et al., 2021), we propose MathFusion, a novel framework that enhances mathematical reasoning by fusing different mathematical problems. The key insight behind MathFusion is that the strategic combination of complementary mathematical instructions can unlock deeper reasoning capabilities. Specifically, by combining two existing problems, MathFusion synthesizes a new math problem that encapsulates the relational and compositional aspects of the original two problems. To achieve this, we introduce three distinct fusion strategies: (1) sequential fusion, which links related problems by chaining them together through shared variables to model solution dependencies; (2) parallel fusion, which integrates analogy problems to enhance conceptual comprehension and generate a novel problem that encapsulates their shared mathematical essence; and (3) conditional fusion, which generates selective problems based on specific context to promote flexible reasoning.
Starting from existing datasets, we first identify pairs of problems that are suitable for fusion. Then we generate new problems by applying these fusion strategies to pairs of mathematical problems that share similar types and contexts. After that, we use strong LLMs to generate corresponding solutions. The resulting dataset, MathFusionQA, is then used to fine-tune LLMs including DeepSeekMath-7B, Mistral-7B, and Llama3-8B.
Experimental results demonstrate that MathFusion enables LLMs to effectively capture the underlying relational structures of mathematical tasks, thereby enhancing their capacity to resolve complex, multi-step problems. Moreover, MathFusion yields considerable improvements in mathematical reasoning accuracy across both in-domain and challenging out-of-domain benchmarks, outperforming traditional single-instruction fine-tuning by 18.0 points in accuracy on average while incorporating only 45K additional synthetic instructions. Further integration with the state-of-the-art (SOTA) data augmentation method DART-Math (Tong et al., 2024) leads to additional improvements, surpassing it by 1.4 points in accuracy on average while utilizing less than one-third of the data employed by DART-Math. This highlights the complementary and orthogonal nature of our approach with existing data augmentation techniques.
2 Related Work
2.1 Individual Data Augmentation for Math
Existing mathematical data augmentation methods primarily focus on two aspects: enhancing existing data and generating new data. Enhancing existing data typically involves modifying the problem or solution. For the problem, strategies include altering the level of complexity/difficulty (Luo et al., 2023), rephrasing the wording (Yu et al., 2024; Li et al., 2024b), and employing backward reasoning (Yu et al., 2024). For the solution, methods such as generating diverse and high-quality mathematical reasoning paths through multiple calls (Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024; Tong et al., 2024), and incorporating reflection (Zhang et al., 2024) are commonly used. Generating new data typically involves creating new mathematical problems based on key mathematical concepts (Tang et al., 2024), seed datasets (Ding et al., 2024), specific example (Li et al., 2024a), and then using strong mathematical models (OpenAI et al., 2023; Shao et al., 2024) to generate corresponding solutions.
These methods, however, focus primarily on individual mathematical problems, overlooking the underlying relationships between different mathematical problems.
2.2 Compositional Data Augmentation
Most data augmentation methods focus on enhancing individual instances, while few consider the relationships between different instances. mixup (Zhang et al., 2018) is an augmentation technique that addresses this gap by generating synthetic training samples through linear interpolations between pairs of input data points and their corresponding labels, which has been shown to be effective across various tasks (Cao et al., 2025; Jin et al., 2024), such as image classification (Zhang et al., 2018; Thulasidasan et al., 2019), text classification (Guo et al., 2019; Zhang et al., 2020), and neural machine translation (Guo et al., 2020; Wu et al., 2021). Mosaic-IT (Li et al., 2024c) is a model-free data augmentation method that concatenates instruction data and trains LLMs with meta-instructions, thereby enhancing performance and reducing training costs. Some works also consider the composition of multiple skills or keypoints. Instruct-SkillMix (Kaur et al., ) extracts core skills for instruction-following and generates new instructions by randomly combining pairs of skills. KPMath (Huang et al., 2024) shares the same idea with Instruct-SkillMix, but focuses on mathematical problems by extracting topics and key points from the problem and generates new problems by combining them.
In contrast to existing works, our approach primarily focuses on fusing mathematical problems and places particular emphasis on the logical coherence of the fusion.
3 MathFusion
The overview of MathFusion is shown in Figure 2. Given two mathematical problems and from the original mathematical training set, MathFusion synthesizes a new mathematical problem by fusing these two problems. A simple example for and is shown in Example 3, and we show the corresponding for three fusion strategies in the following sections. More cases are shown in the Appendix F.
In the following sections, we will first introduce the problem pair construction in Section 3.1, and then introduce the three fusion strategies: sequential fusion in Section 3.2, parallel fusion in Section 3.3, and conditional fusion in Section 3.4. Based on the augmented problem sets generated by these fusion strategies, we present the MathFusionQA dataset in Section 3.5.
3.1 Problem Pair Construction
To construct problem pairs for fusion, for each problem , we need to identify a suitable problem . A straightforward approach is to select a problem that shares the same type and similar context with . Formally, the problem pair set is defined as:
where is a set of problems from the original training set, and is the inner product of the embeddings of and using OpenAI embedding API text-embedding-3-large (OpenAI et al., 2023).
3.2 Sequential Fusion
In mathematical problem-solving, sequential reasoning is a common pattern where the solution of the whole problem is the sequential combination of the solutions of the sub-problems. Sequential fusion constructs a new mathematical problem by establishing solution dependencies between two original problems and through shared variables, where the answer of becomes a prerequisite for solving . Formally, the sequential fusion process and the resulting augmented problem set are defined as:
The answer from solving serves as a part of the input to , thereby creating a chained dependency. A specific example of sequential fusion is shown in Example 3.2. The answer of the (the number of people transported by the boat) is used as the input for (the number of people in the first bus).
3.3 Parallel Fusion
Analogous problems often share common mathematical concepts and essences. Parallel fusion leverages this by synthesizing through the integration of two conceptually analogous problems and , thereby creating a new problem that encapsulates their shared mathematical essence. This approach emphasizes the conceptual relationships between problems rather than their sequential dependencies. The parallel fusion process and the resulting augmented problem set are formally defined as:
where and denote the potentially modified problems from and , respectively, for the fused problem . The function encompasses various operations, such as algebraic composition and the enforcement of constraint satisfiability, to rigorously integrate the underlying mathematical structures. A concrete illustration of parallel fusion is provided in Example 3.3. The total number of people transported by boat and buses over 2 days is asked to be calculated, and the input of (the number of trips made by the boat in one day) is different from that of .
3.4 Conditional Fusion
Context-aware reasoning necessitates the dynamic selection or comparison of solutions based on conditional constraints. Conditional fusion synthesizes by integrating and into a cohesive real-world scenario, where the final solution is derived through contextual comparison or selection of outcomes from and . Formally, the conditional fusion process and the resulting augmented problem set are defined as:
is a comparison function that contrasts and based on predefined logical or contextual rules. A concrete case is shown in Example 3.4, where the final solution is determined by comparing the answers of (the capacity of the boat) and (the capacity of the buses) in a real-world scenario (organizing a lake excursion and a museum trip).
To clarify, the core difference between parallel fusion and conditional fusion is that: parallel fusion combines and to form a novel , where the input of may be different from the original and ; while conditional fusion compares the results of and , the input of is the same as and , and the output is based on the comparison of the results of and .
3.5 MathFusionQA Dataset
After applying the three fusion strategies to and get the augmented problem sets , , and , we use GPT-4o-mini (OpenAI et al., 2023) to generate corresponding solutions for the augmented problems. The resultingaugmented data , , and are combined with the original training set to form the final MathFusionQA dataset as . We use GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) as the original training set separately. We compare our MathFusionQA with other mathematical datasets in Table 1. Though MathFusionQA is overall smaller than other datasets except for RefAug (Zhang et al., 2024), we empirically show that MathFusionQA exhibits strong performance and is more effective than RefAug in Section 4.2. Then we fine-tune LLMs on the MathFusionQA dataset, resulting in MathFusion models. Given that some problems in MathFusionQA may be incomplete or incorrect, we conduct an analysis of problem evaluation and error correction, and present cases of unsuitable fusions in Appendix C.2.
| Dataset | # Samples |
| WizardMath (Luo et al., 2023) | 96K |
| MetaMathQA (Yu et al., 2024) | 395K |
| MMIQC (Liu et al., 2024) | 2294K |
| Orca-Math (Mitra et al., 2024) | 200K |
| Xwin-Math-V1.1 (Li et al., 2024a) | 1440K |
| KPMath-Plus (Huang et al., 2024) | 1576K |
| MathScaleQA (Tang et al., 2024) | 2021K |
| DART-Math-Uniform (Tong et al., 2024) | 591K |
| DART-Math-Hard (Tong et al., 2024) | 585K |
| RefAug (Zhang et al., 2024) | 30K |
| MathFusionQA | 60K |
4 Experiments
4.1 Experimental Setup
Data Synthesis: We use GPT-4o-mini (OpenAI et al., 2023) to fuse the problems and generate the corresponding solutions. The details about generation and corresponding prompts are shown in Appendix B.1 and A).
Training: We conduct standard instruction-tuning on our MathFusionQA. Following DART-Math (Tong et al., 2024), we conduct experiments on two categories of base models: 7B math-specialized base LLM, specifically DeepSeekMath-7B (Shao et al., 2024), and 7-8B general base LLMs, specifically Mistral-7B (Jiang et al., 2023) and Llama3-8B (Dubey et al., 2024). We fine-tune each base model with three fusion strategies—sequential, parallel, and conditional—each of which is the union of GSM8K, MATH, and the augmented set generated by the corresponding fusion strategy. Table 4 shows the statistics of the MathFusionQA collection. All models are trained for 3 epochs for simplicity. More details about the training setup are provided in Appendix B.2.
| Model | # Samples | In-Domain | Out-of-Domain | |||||
| MATH | GSM8K | College | DM | Olympiad | Theorem | AVG | ||
| DeepSeekMath (7B Math-Specialized Base Model) | ||||||||
| DeepSeekMath-7B-RFT | 590K | 53.0 | 88.2 | 41.9 | 60.2 | 19.1 | 27.2 | 48.3 |
| DeepSeekMath-7B-DART-Math | 590K | 53.6 | 86.8 | 40.7 | 61.6 | 21.7 | 32.2 | 49.4 |
| DeepSeekMath-7B-Instruct | 780K | 46.9 | 82.7 | 37.1 | 52.2 | 14.2 | 28.1 | 43.5 |
| DeepSeekMath-7B-MMIQC | 2.3M | 45.3 | 79.0 | 35.3 | 52.9 | 13.0 | 23.4 | 41.5 |
[…]
| DeepSeekMath-7B-RefAug | 30K | 32.1 | 71.2 | 26.0 | 38.4 | 10.1 | 14.4 | 32.0 |
| MathFusion-DSMath-7B (Sequential) | 30K | 49.9 | 76.6 | 38.8 | 64.6 | 21.6 | 22.8 | 45.7 |
| MathFusion-DSMath-7B (Parallel) | 30K | 50.9 | 76.7 | 38.9 | 62.2 | 19.0 | 23.8 | 45.3 |
| MathFusion-DSMath-7B (Conditional) | 30K | 48.5 | 74.6 | 37.0 | 55.2 | 19.3 | 19.0 | 42.3 |
| DeepSeekMath-7B-MetaMath† | 60K | 40.0 | 79.0 | 33.2 | 45.9 | 9.5 | 18.9 | 37.8 |
| DeepSeekMath-7B-MMIQC† | 60K | 26.3 | 60.6 | 19.2 | 41.5 | 10.4 | 6.8 | 27.5 |
| DeepSeekMath-7B-RefAug† | 60K | 33.1 | 71.6 | 26.2 | 35.4 | 10.5 | 14.0 | 31.8 |
| DeepSeekMath-7B-DART-Math† | 60K | 51.4 | 82.9 | 39.1 | 62.8 | 21.0 | 27.4 | 47.4 |
| MathFusion-DSMath-7B | 60K | 53.4 | 77.9 | 39.8 | 65.8 | 23.3 | 24.6 | 47.5 |
| Mistral-7B (7-8B General Base Model) | ||||||||
| Mistral-7B-MetaMath | 400K | 29.8 | 76.5 | 19.3 | 28.0 | 5.9 | 14.0 | 28.9 |
| Mistral-7B-WizardMath-V1.1 | 418K | 32.3 | 80.4 | 23.1 | 38.4 | 7.7 | 16.6 | 33.1 |
| Mistral-7B-RFT | 590K | 38.7 | 82.3 | 24.2 | 35.6 | 8.7 | 16.2 | 34.3 |
| Mistral-7B-DART-Math | 590K | 45.5 | 81.1 | 29.4 | 45.1 | 14.7 | 17.0 | 38.8 |
| Mistral-7B-MathScale | 2.0M | 35.2 | 74.8 | 21.8 | – | – | – | – |
| Mistral-7B-MMIQC | 2.3M | 37.4 | 75.4 | 28.5 | 38.0 | 9.4 | 16.2 | 34.2 |
[…]
| Mistral-7B-RefAug | 30K | 15.1 | 61.1 | 10.4 | 15.4 | 3.1 | 11.0 | 19.4 |
| MathFusion-Mistral-7B (Sequential) | 30K | 32.7 | 73.9 | 18.9 | 29.3 | 9.3 | 15.5 | 29.9 |
| MathFusion-Mistral-7B (Parallel) | 30K | 30.9 | 75.1 | 20.9 | 26.5 | 11.0 | 15.2 | 29.9 |
| MathFusion-Mistral-7B (Conditional) | 30K | 26.3 | 73.0 | 15.6 | 21.4 | 7.3 | 12.8 | 26.1 |
| Mistral-7B-MetaMath† | 60K | 22.7 | 70.8 | 14.1 | 27.2 | 5.0 | 12.2 | 25.3 |
| Mistral-7B-MMIQC† | 60K | 17.3 | 61.4 | 11.1 | 13.5 | 5.0 | 5.9 | 19.0 |
| Mistral-7B-RefAug† | 60K | 17.4 | 63.1 | 12.5 | 18.1 | 3.9 | 11.1 | 21.0 |
| Mistral-7B-DART-Math† | 60K | 34.1 | 77.2 | 23.4 | 36.0 | 8.7 | 18.2 | 32.9 |
| MathFusion-Mistral-7B | 60K | 41.6 | 79.8 | 24.3 | 39.2 | 13.6 | 18.1 | 36.1 |
| Llama3-8B (7-8B General Base Model) | ||||||||
| Llama3-8B-MetaMath | 400K | 32.5 | 77.3 | 20.6 | 35.0 | 5.5 | 13.8 | 30.8 |
| Llama3-8B-RFT | 590K | 39.7 | 81.7 | 23.9 | 41.7 | 9.3 | 14.9 | 35.2 |
| Llama3-8B-MMIQC | 2.3M | 39.5 | 77.6 | 29.5 | 41.0 | 9.6 | 16.2 | 35.6 |
| Llama3-8B-DART-Math | 590K | 46.6 | 81.1 | 28.8 | 48.0 | 14.5 | 19.4 | 39.7 |
[…]
| Llama3-8B-RefAug | 30K | 20.8 | 67.3 | 15.7 | 25.9 | 4.7 | 13.6 | 24.7 |
| MathFusion-Llama3-8B (Sequential) | 30K | 38.8 | 77.9 | 25.1 | 42.0 | 12.6 | 17.0 | 35.6 |
| MathFusion-Llama3-8B (Parallel) | 30K | 38.1 | 75.4 | 25.5 | 41.9 | 11.9 | 18.9 | 35.3 |
| MathFusion-Llama3-8B (Conditional) | 30K | 34.7 | 76.9 | 21.2 | 27.4 | 11.9 | 15.5 | 31.3 |
| Llama3-8B-MetaMath† | 60K | 28.7 | 78.5 | 19.7 | 31.3 | 5.3 | 16.1 | 29.9 |
| Llama3-8B-MMIQC† | 60K | 24.4 | 69.7 | 13.4 | 30.9 | 5.2 | 10.6 | 25.7 |
| Llama3-8B-RefAug† | 60K | 20.3 | 68.6 | 15.5 | 29.1 | 5.5 | 13.0 | 25.3 |
| Llama3-8B-DART-Math† | 60K | 39.6 | 82.2 | 27.9 | 39.9 | 12.9 | 22.9 | 37.6 |
| MathFusion-Llama3-8B | 60K | 46.5 | 79.2 | 27.9 | 43.4 | 17.2 | 20.0 | 39.0 |
Evaluation: Following DART-Math (Tong et al., 2024), we evaluate the models on two in-domain (ID) benchmarks: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), as our MathFusionQA dataset is built upon these two datasets. For out-of-domain (OOD) evaluation, we use the CollegeMath (Tang et al., 2024), DeepMind-Mathematics (Saxton et al., 2019), OlympiadBench-Math (He et al., 2024), and TheoremQA (Chen et al., 2023) benchmarks. We use greedy decoding to generate solutions for the problems in test sets. We report the accuracy in 0-shot setting for all models following Tong et al. (2024). More details about the evaluation setup and benchmarks are provided in the Appendix B.3.
Baselines: We mainly compare our MathFusion models with mathematical instruction-based models, which can be categorized into three groups: (1) Previous top-performing models, including MetaMath (Yu et al., 2024), WizardMath (Luo et al., 2023), RFT (rejection sampling fine-tuning) (Yuan et al., 2023; Tong et al., 2024), MMIQC (Liu et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (Shao et al., 2024), RefAug (Zhang et al., 2024), and DART-Math (Tong et al., 2024) (we report the Prop2Diff version); (2) Models instruction-tuned on the combination of GSM8K and MATH datasets (noted as ‘standard’ setting); (3) Models instruction-tuned on the sampled 60K version of previous top-performing methods to further evaluate the data efficiency of different mathematical data augmentation methods. Details about the sampling method are introduced in Appendix B.2.
4.2 Main Results
The main results are shown in Table 2. We summarize several key findings as follows:
Finding 1: Three fusion strategies consistently enhance the model performance. For all three fusion strategies-sequential, parallel, and conditional fusion—the MathFusion models consistently surpass the standard settings across all base models and evaluation benchmarks. Specifically, on MATH and GSM8K test sets, using Llama3-8B as the base model, MathFusion (sequential) achieves 21.3 and 12.5 accuracy improvement; MathFusion (parallel) achieves 20.6 and 10.0 accuracy improvement; and MathFusion (conditional) achieves 18.0 and 11.9 accuracy improvement, respectively, compared to the standard setting. For four OOD benchmarks, the single fusion strategy also outperforms the standard setting, with a 9.9 accuracy improvement on average. These improvements demonstrate the effectiveness of the three fusion strategies in enhancing both the ID and OOD generalization performance of the models.
Finding 2: Among three fusion strategies, sequential fusion and parallel fusion generally perform better than conditional fusion. A possible reason is that the conditional fusion requires no modification of input structures or problem dependencies, merely performing a direct comparison or selection between the solutions of two independent problems without necessitating additional mathematical transformations or reformulations. We further investigate the difficulty of the problems generated by the three fusion strategies in Section 5.1.
Finding 3: Combination of three fusion strategies further improves performance. As the three fusion strategies capture different aspects of the problem fusion, we further investigate the performance of the combined fusion strategies. From Table 2, we observe that the combined fusion strategies consistently outperform each single fusion strategy, indicating that the combination of three fusion strategies can further enhance the model’s mathematical ability. Additionally, the weaker the performance of the base model, the more enhancements the combined fusion strategies can bring. Specifically, the combined fusion strategies achieve an average accuracy improvement of 3.1 points on DeepSeekMath-7B, 4.9 points on Llama3-8B, and 7.5 points on Mistral-7B across all benchmarks.
Finding 4: Compared with previous top-performing baselines, MathFusion models yields competitive performance and high data efficiency. For each single fusion strategy, MathFusion models outperform RefAug, which has the same data size as MathFusion, on all benchmarks. After combining the three fusion strategies, MathFusion outperforms previous top-performing baselines like MetaMath and DART-Math on average under the same data size setting. Specifically, MathFusion yields consistently better performance on MATH, DeepMind-Mathematics, and OlympiadBench-Math benchmarks. These results demonstrate the high data efficiency and generalization ability of MathFusion. MathFusion maintains also competitive efficacy compared to top-performing models in the full-data regime, exhibiting only a marginal average performance drop on Llama3-8B and DeepSeekMath-7B.
4.3 Ablation Study
| Method | Sequential | Parallel | Conditional | MATH | GSM8K |
| Standard | ✗ | ✗ | ✗ | 17.5 | 65.4 |
| MathFusion | ✗ | ✓ | ✓ | 42.6 | 78.2 |
| ✓ | ✗ | ✓ | 43.0 | 76.9 | |
| ✓ | ✓ | ✗ | 43.6 | 79.2 | |
| ✓ | ✓ | ✓ | 45.6 | 79.9 |
We further conduct an ablation study to investigate the contribution of each fusion strategy to the overall performance of combined fusion. The results over Llama3-8B on MATH and GSM8K are shown in Table 3, from which we observe that each fusion strategy contributes to the overall performance, with conditional fusion showing the least contribution, which aligns with Section 4.2.
5 Analysis
5.1 Difficulty Analysis
In this section, we explore why the three fusion strategies effectively enhance the model’s performance. To achieve this, we evaluate both the perplexity (PPL) and instruction following difficulty (IFD) (Li et al., 2024d) for the original and fused data. We use Mathstral-7B (team, 2024), a model built upon Mistral-7B (Jiang et al., 2023) and specifically fine-tuned for mathematical reasoning, to ensure our analysis relies on a model specifically designed for mathematical tasks. Specifically, we denote the unconditioned PPL as , the conditioned PPL as , and , where is the problem and is the solution. The results are shown in Figure 3 and 3, from which we can see: (1) The PPL of the solution of the fused problems is significantly lower than that of the original problems. As analyzed in Yu et al. (2024), this may be due to the easy-to-learn nature of the generated solutions. (2) The IFD of the fused data is significantly higher than that of the original data, indicating that the fused data is more difficult to learn in the context of the problem. (3) The IFD of the MATH datasets, both the original or fused version, are higher than that of the GSM8K, consistent with the fact that MATH is generally more difficult than GSM8K.
5.2 Relationship between Augmented Data Size and Performance
We study the performance scaling behavior of the MathFusion on different sizes of augmented data on Llama3-8B. We select MATH as the original training set and gradually increase the size of the augmented fusion data from 0 to 22.5K, with a step size of 2.5K. The results on MATH and four OOD benchmarks are shown in Figure 3. We observe that the performance of the MathFusion models exhibits an approximate logarithmic growth with respect to the amount of augmented data, which is consistent with the findings in (Li et al., 2024b). Additionally, the augmented fusion data from MATH dataset can also generalize better to the OOD benchmarks as the size of the augmented data increases. In summary, the MathFusion shows consistent performance improvement with different sizes of augmented data.
5.3 Combination with Other Datasets
We further investigate the performance of MathFusion when combined with other data augmentation methods. Specifically, we downsample 30K-180K data from DART-Math-Hard (Tong et al., 2024), which is the SOTA method for mathematical data augmentation with 590K data. We combine the downsampled DART-Math-Hard with our MathFusionQA dataset and fine-tune Llama3-8B models on the combined dataset. The results are presented in Table 4. As the size of sampled data increases, the average performance of the models also increases, and reaches the peak when the size of the sampled data is 120K. Notably, by only using 90K data sampled from DART-Math-Hard (i.e., 150K samples in total), the resulting model achieves better performance than both DART-Math and MathFusion, yields SOTA average performance. These results show the potential of combining MathFusion with other data augmentation methods to further enhance the model’s performance. We think that the enhancement arises from the complementary and orthogonal nature of the two methods: our MathFusion emphasizes fusing mathematical problems to generate more challenging and diverse problems, while DART-Math focuses on existing difficult problems and primarily generates additional solutions for them.
5.4 Diversity Analysis
To further investigate the effectiveness of the MathFusion in enhancing the data diversity, we visualize the problem embeddings of the GSM8K and MATH datasets generated by GPT-4o-mini using t-SNE (Van der Maaten and Hinton, 2008). The results are shown in Figure 4 and 4. We can observe that the MathFusion augmented problems are more evenly distributed in the embedding space, thereby enriching the diversity of the training examples and mitigating the risk of model overfitting.
6 Conclusion
In this paper, we focus on the fusion of mathematical problems. We propose a novel mathematical data augmentation method, MathFusion, which comprises three distinct fusion strategies—sequential fusion, parallel fusion, and conditional fusion—designed to synthesize augmented mathematical problems. Leveraging these fusion strategies, we construct the MathFusionQA dataset, which is subsequently employed to fine-tune LLMs. Extensive experiments on three base models and six benchmarks show that MathFusion exhibits robust performance in both the in-domain and out-of-domain benchmarks while maintaining high data efficiency.
Limitations
We utilize strong GPT-4o-mini to generate fused problems and solutions, but the generated problems or solutions may still contain errors or ambiguities, which are hard to detect and verify. The quality of the generated problems and solutions is limited by the capabilities of the teacher LLM. We mainly explore the effectiveness of the three fusion strategies on problem pairs that are constructed by embedding similarity. The fusion of three or more problems and more effective ways to find similar problems, remain unexplored.
Appendix A Prompts
We show the prompts used for Sequential Fusion in Prompt E, Parallel Fusion in Prompt E, and Conditional Fusion in Prompt E. We also provide the problem evaluation prompts in Prompt E, which is partially derived from WizardMath (Luo et al., 2023). We use LangGPT (Wang et al., 2024b) to format prompts in Markdown and polish them.
Appendix B General Settings
B.1 Data Synthesis
We synthesize the augmented data, both the fusion process and the generation of the corresponding solutions, using GPT-4o-mini(gpt-4o-mini-2024-07-18) (OpenAI et al., 2023). We set the temperature to 0.7 and the maximum length of generation to 4096. The statistics of the generated data, as well as the base GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets, are shown in Table 4.
| Dataset | GSM8K | MATH | Total |
| Standard | 7.5K | 7.5K | 15K |
| MathFusionQA (Sequential) | 15K | 15K | 30K |
| MathFusionQA (Parallel) | 15K | 15K | 30K |
| MathFusionQA (Conditional) | 15K | 15K | 30K |
| MathFusionQA | 30K | 30K | 60K |
B.2 Training
We use LLaMA-Factory (Zheng et al., 2024) to fine-tune the models. All models, including our own reproductions of baselines, are fine-tuned for 3 epochs with a batch size of 128 on 8xNVIDIA A100 GPU. The peak learning rate is 5e-6 with a linear warm-up for the first 3% of the training steps, followed by cosine decay. The maximum sequence length is set to 4096.
In Table 2, we reproduce the results of the baselines with 60K data. For MetaMath (Yu et al., 2024), MMIQC (Liu et al., 2024), and DART-Math (Tong et al., 2024), we directly downsample 60K data from the original datasets randomly. For RefAug (Zhang et al., 2024), the original training set only contains 30K data, with 15K from GSM8K and MATH, and 15K from the augmented reflection data. To upsample the RefAug dataset to 60K, we re-generate the reflection data two times using GPT-4o-mini with the original prompts (Zhang et al., 2024), thus obtaining an additional 30K data and forming the 60K dataset.
B.3 Evaluation
We compare MathFusion models with baselines on the following six benchmarks:
GSM8K (Cobbe et al., 2021) dataset includes 8,792 high-quality grade school math word problems, with 7,473 for training and 1,319 for testing. Each problem in GSM8K requires between 2 and 8 steps to solve.
MATH (Hendrycks et al., 2021) dataset is composed of 12,500 problems from high school math competitions, with 7,500 for training and 5,000 for testing. Problems in MATH are categorized into 7 types (Prealgebra, Intermediate Algebra, Algebra, Precalculus, Geometry, Counting & Probability, and Number Theory) and 5 difficulty levels.
CollegeMath (Tang et al., 2024) test set contains 2,818 college-level problems, which are curated from 9 college-level mathematics textbooks, covering 7 key mathematical disciplines: Algebra, Precalculus, Calculus, VectorCalculus, Probability, LinearAlgebra, and Differential Equations.
DeepMind-Mathematics (Saxton et al., 2019) test set consists of 1,000 problems covering a wide range of mathematical reasoning tasks spanning algebra, arithmetic, calculus, and probability designed to evaluate the mathematical reasoning abilities of models.
OlympiadBench-Math (He et al., 2024) benchmark including 675 Olympiad-level mathematical problems, and we only use the text-only English subset of Olympiad-Bench.
TheoremQA (Chen et al., 2023) is a novel theorem-driven question-answering benchmark containing 800 problems based on 350 theorems. It is designed to evaluate LLM’s ability to apply domain-specific theorems across fields such as Mathematics, Physics, Electrical Engineering, Computer Science, and Finance.
B.4 Templates
[…]
Appendix C Analysis of Fused Problems
The embedding search naturally ensures a high degree of contextual similarity. In the following sections, we analyze the fused problems in terms of problem types and errors.
C.1 Fused Probelm Types
Regarding problem types, in the GSM8K (Cobbe et al., 2021) dataset, all problems are simple algebra questions. For the MATH dataset, we find that 83% of the problem pairs belong to the same category, further validating the feasibility of the embedding search. We plot the distribution of combination types of problems in MATH in Figure 5.
| Model | In-Domain | Out-of-Domain | |||||
| MATH | GSM8K | College | DM | Olympiad | Theorem | AVG | |
| Standard #1 | 17.4 | 63.1 | 12.1 | 23.1 | 3.7 | 9.6 | 21.5 |
| Standard #2 | 17.6 | 63.7 | 12.6 | 20.6 | 4.3 | 8.9 | 21.3 |
| Standard #3 | 17.5 | 65.4 | 12.9 | 21.6 | 4.7 | 10.9 | 22.2 |
| Standard (Avg.) | 17.50.1 | 64.11.2 | 12.50.4 | 21.81.3 | 4.20.5 | 9.81.0 | 21.70.5 |
| MathFusion #1 | 45.6 | 79.9 | 27.1 | 44.4 | 17.2 | 19.5 | 39.0 |
| MathFusion #2 | 45.3 | 79.8 | 27.5 | 45.4 | 17.0 | 19.4 | 39.1 |
| MathFusion #3 | 46.5 | 79.2 | 27.9 | 43.4 | 17.2 | 20.0 | 39.0 |
| MathFusion(Avg.) | 45.80.6 | 79.60.4 | 27.50.4 | 44.41.0 | 17.10.1 | 19.60.3 | 39.00.1 |
| Model | # Samples | In-Domain | Out-of-Domain | |||||
| MATH | GSM8K | College | DM | Olympiad | Theorem | AVG | ||
| Standard | 15K | 17.5 | 65.4 | 12.9 | 21.6 | 4.7 | 10.9 | 22.2 |
| Standard + GPT Rewritten | 30K | 22.8 | 75.4 | 11.8 | 15.7 | 5.5 | 9.6 | 23.5 |
| MathFusion (Sequential) | 30K | 38.8 | 77.9 | 25.1 | 42.0 | 12.6 | 17.0 | 35.6 |
| MathFusion (Parallel) | 30K | 38.1 | 75.4 | 25.5 | 41.9 | 11.9 | 18.9 | 35.3 |
| MathFusion (Conditional) | 30K | 34.7 | 76.9 | 21.2 | 27.4 | 11.9 | 15.5 | 31.3 |
| MathFusion | 60K | 46.5 | 79.2 | 27.9 | 43.4 | 17.2 | 20.0 | 39.0 |
C.2 Fused Error Analysis
In practice, we find that some fused problems are unreasonable or ambiguous, which are shown in Section F. The reason may be that some problems are not suitable for fusion or the limited capacity of the model for generating fused problems. To verify the correctness of the fused problems and their influence on the model’s performance, we conduct an error analysis on the fused problems. Specifically, borrowing the idea from rejection sampling (Yuan et al., 2023), we use GPT-4o-mini to verify the correctness and completeness of the fused problems. The corresponding evaluation prompt is shown in Section A. For each identified unreasonable problem, we adjust the temperature to 1.0 to enhance the diversity of generation, and re-generate the problems five times using the corresponding fusion strategy. If none of the five generated problems is reasonable, we consider the fusion to be unreasonable and discard it. Finally, 5.6% of the fused problems are identified as unreasonable, and the remaining reasonable problems are added to the dataset. The average performance of Llama3-8B fine-tuned only on the filtered MathFusionQA is 39.1, which is similar to the performance of the model fine-tuned on the original MathFusionQA (39.0), indicating that the unreasonable problems have little impact on the model’s performance.
Appendix D Effect of Teacher Model
In MathFusion, we use GPT-4o-mini (OpenAI et al., 2023) as the teacher model to generate the solutions for the fused problems. To validate the performance improvement of MathFusion is not merely due to the stronger teacher model, we conduct an ablation study, where we use GPT-4o-mini to rewrite the solutions from the original training set. Then we fine-tune the Llama3-8B model on the original training set and the rewritten solutions. The results are shown in Table 6. We can see that the performance of the model fine-tuned on the rewritten solutions is better than the Standard setting, especially on the MATH and GSM8K datasets. However, the average improvement is only 1.3 points. Meanwhile, each fusion strategy of MathFusion still outperforms the rewritten solution by a large margin, indicating that the performance improvement of MathFusion mainly comes from the fusion of problems rather than the stronger teacher model.
Appendix E Significant Test
We conduct error analysis on MathFusion on Llama3-8B model to verify the consistent performance improvement of our MathFusionQA. Specifically, we fine-tune the Llama3-8B model on the original training sets (Standard setting), and the combined fusion strategies, respectively. The results are shown in Table 5. We can see that the MathFusion models consistently outperform the standard setting across all benchmarks. We also conduct statistical significance tests using the paired t-test, and results show that the performance improvement of MathFusion is statistically significant () on all benchmarks.
Appendix F More Cases
More cases, including the original problems and , the fused problem , are shown below. Specifically, we show three reasonable cases in Case F, Case F, and Case F, and three unreasonable cases in Case F, Case F, and Case F.

### v2

@@
-MathFusion: Enhancing Mathematic Problem-solving of LLM through Instruction Fusion
+MathFusion: Enhancing Mathematical Problem-solving of LLM through Instruction Fusion
 Abstract
@@
 Experimental results demonstrate that MathFusion achieves substantial improvements in mathematical reasoning while maintaining high data efficiency, boosting performance by 18.0 points in accuracy across diverse benchmarks while requiring only 45K additional synthetic instructions, representing a substantial improvement over traditional single-instruction approaches.
-Our datasets, models, and code are publicly available at https://github.com/QizhiPei/mathfusion.
-MathFusion: Enhancing Mathematic Problem-solving of LLM
+MathFusion: Enhancing Mathematical Problem-solving of LLM
 through Instruction Fusion
-Qizhi Pei1, Lijun Wu2∗, Zhuoshi Pan3, Yu Li2, Honglin Lin2, Chenlin Ming4, Xin Gao2, Conghui He2∗, Rui Yan1,5††thanks: Corresponding authors: Lijun Wu (wulijun@pjlab.org.cn), Conghui He (heconghui@pjlab.org.cn), and Rui Yan (ruiyan@ruc.edu.cn).
-1Gaoling School of Artificial Intelligence, Renmin University of China 2Shanghai AI Laboratory 3Tsinghua University 4Shanghai Jiao Tong University 5School of Computer Science, Wuhan University {qizhipei,ruiyan}@ruc.edu.cn {wulijun,heconghui}@pjlab.org.cn
+Qizhi Pei1,2,
+Lijun Wu2∗,
+Zhuoshi Pan2,3,
+Yu Li2,
+Honglin Lin2,
+Chenlin Ming2,4,
+Xin Gao2,
+Conghui He2∗,
+Rui Yan1,5,6††thanks: Corresponding authors: Lijun Wu (wulijun@pjlab.org.cn), Conghui He (heconghui@pjlab.org.cn), and Rui Yan (ruiyan@ruc.edu.cn).
+1Gaoling School of Artificial Intelligence, Renmin University of China
+2Shanghai Artificial Intelligence Laboratory 3Tsinghua University 4Shanghai Jiao Tong University
+5Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE
+6School of Artificial Intelligence, Wuhan University
+{qizhipei,ruiyan}@ruc.edu.cn {wulijun,heconghui}@pjlab.org.cn
+https://huggingface.co/collections/QizhiPei/MathFusion
+https://github.com/QizhiPei/MathFusion
 1 Introduction
@@
 Specialized mathematical LLMs have emerged to address the unique challenges of solving complex mathematical problems (Yang et al., 2024; Shao et al., 2024; Ying et al., 2024; team, 2024).
-Current approaches to enhance mathematical reasoning primarily focus on four paradigms: continued pre-training with math corpora (Yang et al., 2024; Shao et al., 2024), reinforcement learning (RL) from human or automated feedback (Luo et al., 2023; Lu et al., 2024), test-time compute scaling (Wang et al., 2024a; Kang et al., 2024; Guan et al., 2025; Xi et al., 2024), and supervised fine-tuning (SFT) using problem-solution pairs (Tang et al., 2024; Tong et al., 2024).
+Current approaches to enhance mathematical reasoning primarily focus on four paradigms: continued pre-training with math corpora (Yang et al., 2024; Shao et al., 2024), reinforcement learning (RL) from human or automated feedback (Luo et al., 2023; Lu et al., 2024), test-time compute scaling (Wang et al., 2024a; Kang et al., 2024; Guan et al., 2025; Xi et al., 2024), and supervised fine-tuning (SFT) using problem-solution pairs (Tang et al., 2024; Tong et al., 2024; Lin et al., 2025).
 Among these, SFT is the most widely adopted paradigm (Setlur et al., 2024) due to its simplicity.
@@
 Specifically, by combining two existing problems, MathFusion synthesizes a new math problem that encapsulates the relational and compositional aspects of the original two problems.
-To achieve this, we introduce three distinct fusion strategies: (1) sequential fusion, which links related problems by chaining them together through shared variables to model solution dependencies; (2) parallel fusion, which integrates analogy problems to enhance conceptual comprehension and generate a novel problem that encapsulates their shared mathematical essence; and (3) conditional fusion, which generates selective problems based on specific context to promote flexible reasoning.
+To achieve this, we introduce three distinct fusion strategies: (1) sequential fusion, which links related problems by chaining them together through shared variables to model solution dependencies; (2) parallel fusion, which integrates analogy problems to enhance conceptual comprehension and generates a novel problem that encapsulates their shared mathematical essence; and (3) conditional fusion, which generates selective problems based on specific context to promote flexible reasoning.
 Starting from existing datasets, we first identify pairs of problems that are suitable for fusion.
@@
 Moreover, MathFusion yields considerable improvements in mathematical reasoning accuracy across both in-domain and challenging out-of-domain benchmarks, outperforming traditional single-instruction fine-tuning by 18.0 points in accuracy on average while incorporating only 45K additional synthetic instructions.
-Further integration with the state-of-the-art (SOTA) data augmentation method DART-Math (Tong et al., 2024) leads to additional improvements, surpassing it by 1.4 points in accuracy on average while utilizing less than one-third of the data employed by DART-Math.
-This highlights the complementary and orthogonal nature of our approach with existing data augmentation techniques.
+Further integration with the state-of-the-art (SOTA) data augmentation method DART-Math (Tong et al., 2024) and scaling MathFusion to larger size lead to additional improvements, surpassing DART-Math in accuracy on average while utilizing less than one-third of its data.
+This highlights the complementary nature and scalability of our approach.
 2 Related Work
@@
 For the problem, strategies include altering the level of complexity/difficulty (Luo et al., 2023), rephrasing the wording (Yu et al., 2024; Li et al., 2024b), and employing backward reasoning (Yu et al., 2024).
-For the solution, methods such as generating diverse and high-quality mathematical reasoning paths through multiple calls (Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024; Tong et al., 2024), and incorporating reflection (Zhang et al., 2024) are commonly used.
+For the solution, methods such as generating diverse and high-quality mathematical reasoning paths through multiple calls (Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024; Tong et al., 2024), and incorporating reflection (Zhang et al., 2024; Pan et al., 2025) are commonly used.
 Generating new data typically involves creating new mathematical problems based on key mathematical concepts (Tang et al., 2024), seed datasets (Ding et al., 2024), specific example (Li et al., 2024a), and then using strong mathematical models (OpenAI et al., 2023; Shao et al., 2024) to generate corresponding solutions.
-These methods, however, focus primarily on individual mathematical problems, overlooking the underlying relationships between different mathematical problems.
+These methods, however, focus primarily on individual mathematical problems, overlooking the underlying relationships between different problems.
 2.2 Compositional Data Augmentation
@@
 A simple example for and is shown in Example 3, and we show the corresponding for three fusion strategies in the following sections.
-More cases are shown in the Appendix F.
+More cases are shown in the Appendix G.
 In the following sections, we will first introduce the problem pair construction in Section 3.1, and then introduce the three fusion strategies: sequential fusion in Section 3.2, parallel fusion in Section 3.3, and conditional fusion in Section 3.4.
@@
 A specific example of sequential fusion is shown in Example 3.2.
-The answer of the (the number of people transported by the boat) is used as the input for (the number of people in the first bus).
+The answer of (the number of people transported by the boat) is used as the input for (the number of people in the first bus).
 3.3 Parallel Fusion
@@
 Following DART-Math (Tong et al., 2024), we conduct experiments on two categories of base models: 7B math-specialized base LLM, specifically DeepSeekMath-7B (Shao et al., 2024), and 7-8B general base LLMs, specifically Mistral-7B (Jiang et al., 2023) and Llama3-8B (Dubey et al., 2024).
-We fine-tune each base model with three fusion strategies—sequential, parallel, and conditional—each of which is the union of GSM8K, MATH, and the augmented set generated by the corresponding fusion strategy.
+Each base model is fine-tuned using three distinct fusion strategies: sequential, parallel, and conditional.
+For each strategy, the fine-tuning dataset (30K samples) comprises the union of the GSM8K, MATH datasets, and an augmented set generated by the respective fusion strategy.
+The MathFusionQA dataset (60K samples) is formed by the union of all these sub-datasets.
 Table 4 shows the statistics of the MathFusionQA collection.
-All models are trained for 3 epochs for simplicity.
+To demonstrate the scaling ability of MathFusion, we also enlarge MathFusionQA dataset by including top-2 to top-4 nearest neighbors, resulting in a total of 15K + 4 × (3 × 15K) = 195K examples.
+All models are trained for 3 epochs.
 More details about the training setup are provided in Appendix B.2.
@@
 | DeepSeekMath-7B-MMIQC | 2.3M | 45.3 | 79.0 | 35.3 | 52.9 | 13.0 | 23.4 | 41.5 |
+| MathFusion-DSMath-7B | 195K | 58.2 | 79.5 | 40.3 | 69.1 | 25.5 | 27.0 | 49.9 |
 […]
@@
 More details about the evaluation setup and benchmarks are provided in the Appendix B.3.
-Baselines: We mainly compare our MathFusion models with mathematical instruction-based models, which can be categorized into three groups: (1) Previous top-performing models, including MetaMath (Yu et al., 2024), WizardMath (Luo et al., 2023), RFT (rejection sampling fine-tuning) (Yuan et al., 2023; Tong et al., 2024), MMIQC (Liu et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (Shao et al., 2024), RefAug (Zhang et al., 2024), and DART-Math (Tong et al., 2024) (we report the Prop2Diff version); (2) Models instruction-tuned on the combination of GSM8K and MATH datasets (noted as ‘standard’ setting); (3) Models instruction-tuned on the sampled 60K version of previous top-performing methods to further evaluate the data efficiency of different mathematical data augmentation methods.
+Baselines: We mainly compare our MathFusion models with mathematical instruction-based models, which can be categorized into three groups: (1) Previous top-performing models, including MetaMath (Yu et al., 2024), WizardMath (Luo et al., 2023), RFT (rejection sampling fine-tuning from DART-Math) (Yuan et al., 2023; Tong et al., 2024), MMIQC (Liu et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (Shao et al., 2024), RefAug (Zhang et al., 2024), and DART-Math (Tong et al., 2024) (we report the Prop2Diff version as it generally performs better than the Uniform version); (2) Models instruction-tuned on the combination of GSM8K and MATH datasets (noted as “Standard” setting); (3) Models instruction-tuned on the sampled 60K version of previous top-performing methods to further evaluate the data efficiency of different mathematical data augmentation methods.
 Details about the sampling method are introduced in Appendix B.2.
@@
 MathFusion maintains also competitive efficacy compared to top-performing models in the full-data regime, exhibiting only a marginal average performance drop on Llama3-8B and DeepSeekMath-7B.
+Finding 5: MathFusion exhibits strong scalability and outperforms larger-scale baselines with fewer samples.
+The results on DeepSeekMath-7B (the strongest base model for math) in Table 2 reveal that scaling MathFusion from 60K to 195K samples leads to consistent performance gains across all evaluation benchmarks.
+Notably, MathFusion-DSMath-7B (195K) surpasses DeepSeekMath-7B-DART-Math (590K) in average accuracy (49.9 vs.
+49.4), despite using only one-third of the training data.
+This illustrates the high scalability and data efficiency of MathFusion.
+The substantial performance gains on challenging benchmarks such as MATH (+4.6), DeepMind-Mathematics (+7.5), and OlympiadBench-Math (+3.8) underscore the method’s capability to generalize well under increased data volume.
+These findings demonstrate that MathFusion benefits substantially from larger synthetic training sets and can outperform significantly larger instruction-tuned models with less data.
 4.3 Ablation Study
@@
 The results over Llama3-8B on MATH and GSM8K are shown in Table 3, from which we observe that each fusion strategy contributes to the overall performance, with conditional fusion showing the least contribution, which aligns with Section 4.2.
+We further ablate on choice of teacher model for solution generation in Appendix D.
 5 Analysis
+We analyze the difficulty of the fused problem in Section 5.1, the relationship between augmented data size and performance in Section 5.2, the combination of MathFusionQA with other datasets in Section 5.3, and the diversity of fused problem in Section 5.4.
 5.1 Difficulty Analysis
@@
 We combine the downsampled DART-Math-Hard with our MathFusionQA dataset and fine-tune Llama3-8B models on the combined dataset.
-The results are presented in Table 4.
+The results are presented in Figure 4.
 As the size of sampled data increases, the average performance of the models also increases, and reaches the peak when the size of the sampled data is 120K.
@@
 Limitations
-We utilize strong GPT-4o-mini to generate fused problems and solutions, but the generated problems or solutions may still contain errors or ambiguities, which are hard to detect and verify.
+We utilize GPT-4o-mini to generate fused problems and solutions, but the generated problems or solutions may still contain errors or ambiguities, which are hard to detect and verify.
 The quality of the generated problems and solutions is limited by the capabilities of the teacher LLM.
+Stronger teacher model, like DeepSeek-R1 and Qwen3, are underexplored.
 We mainly explore the effectiveness of the three fusion strategies on problem pairs that are constructed by embedding similarity.
-The fusion of three or more problems and more effective ways to find similar problems, remain unexplored.
+The fusion of three or more problems and more effective ways to find similar problems, remain underexplored.
+The released MathFusionQA dataset currently contains only 60K examples, and scaling to millions of examples remains underexplored.
+Acknowledgements
+This work is supported by Beijing Outstanding Young Scientist Program NO.
+BJJWZYJH012019100020098 and by National Key R&D Program of China (2022ZD0160201).
+This work is also supported by the Public Computing Cloud, Renmin University of China and by fund for building world-class universities (disciplines) of Renmin University of China.
+Qizhi Pei is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China.
+Qizhi Pei is an intern at Shanghai Artificial Intelligence Laboratory.
 Appendix A Prompts
-We show the prompts used for Sequential Fusion in Prompt E, Parallel Fusion in Prompt E, and Conditional Fusion in Prompt E.
-We also provide the problem evaluation prompts in Prompt E, which is partially derived from WizardMath (Luo et al., 2023).
+We show the prompts used for Sequential Fusion in Prompt F, Parallel Fusion in Prompt F, and Conditional Fusion in Prompt F.
+We also provide the problem evaluation prompts in Prompt F, which is partially derived from WizardMath (Luo et al., 2023).
 We use LangGPT (Wang et al., 2024b) to format prompts in Markdown and polish them.
@@
 | MATH | GSM8K | College | DM | Olympiad | Theorem | AVG | ||
-| Standard | 15K | 17.5 | 65.4 | 12.9 | 21.6 | 4.7 | 10.9 | 22.2 |
-| Standard + GPT Rewritten | 30K | 22.8 | 75.4 | 11.8 | 15.7 | 5.5 | 9.6 | 23.5 |
-| MathFusion (Sequential) | 30K | 38.8 | 77.9 | 25.1 | 42.0 | 12.6 | 17.0 | 35.6 |
-| MathFusion (Parallel) | 30K | 38.1 | 75.4 | 25.5 | 41.9 | 11.9 | 18.9 | 35.3 |
-| MathFusion (Conditional) | 30K | 34.7 | 76.9 | 21.2 | 27.4 | 11.9 | 15.5 | 31.3 |
-| MathFusion | 60K | 46.5 | 79.2 | 27.9 | 43.4 | 17.2 | 20.0 | 39.0 |
+| Standard | 15K | 12.4 | 60.3 | 8.4 | 17.0 | 2.2 | 7.6 | 18.0 |
+| GPT Rewritten | 15K | 20.1 | 70.3 | 9.1 | 13.9 | 2.8 | 8.9 | 20.9 |
+| Mosaic-IT | 15K | 11.7 | 40.9 | 7.4 | 9.2 | 2.7 | 9.9 | 13.6 |
+| Mosaic-IT + Original GSM8K and MATH | 30K | 11.0 | 54.7 | 6.9 | 9.8 | 1.9 | 9.5 | 15.6 |
+| DART-Math | 60K | 34.1 | 77.2 | 23.4 | 36.0 | 8.7 | 18.2 | 32.9 |
+| MathFusion (Sequential) | 30K | 32.7 | 73.9 | 18.9 | 29.3 | 9.3 | 15.5 | 29.9 |
+| MathFusion (Parallel) | 30K | 30.9 | 75.1 | 20.9 | 26.5 | 11.0 | 15.2 | 29.9 |
+| MathFusion (Conditional) | 30K | 26.3 | 73.0 | 15.6 | 21.4 | 7.3 | 12.8 | 26.1 |
+| MathFusion (DeepSeekMath-7B-RL) | 60K | 42.0 | 78.1 | 24.0 | 36.5 | 13.0 | 13.8 | 34.6 |
+| MathFusion | 60K | 41.6 | 79.8 | 24.3 | 39.2 | 13.6 | 18.1 | 36.1 |
 C.2 Fused Error Analysis
-In practice, we find that some fused problems are unreasonable or ambiguous, which are shown in Section F.
+In practice, we find that some fused problems are unreasonable or ambiguous, which are shown in Section G.
 The reason may be that some problems are not suitable for fusion or the limited capacity of the model for generating fused problems.
@@
 The average performance of Llama3-8B fine-tuned only on the filtered MathFusionQA is 39.1, which is similar to the performance of the model fine-tuned on the original MathFusionQA (39.0), indicating that the unreasonable problems have little impact on the model’s performance.
+This result aligns with the findings in OpenMathInstruct-2 (Toshniwal et al., ), indicating models exhibit some robustness to low-quality data in SFT.
 Appendix D Effect of Teacher Model
 In MathFusion, we use GPT-4o-mini (OpenAI et al., 2023) as the teacher model to generate the solutions for the fused problems.
-To validate the performance improvement of MathFusion is not merely due to the stronger teacher model, we conduct an ablation study, where we use GPT-4o-mini to rewrite the solutions from the original training set.
-Then we fine-tune the Llama3-8B model on the original training set and the rewritten solutions.
+To validate the performance improvement of MathFusion is not merely due to the stronger teacher model, we conduct two ablation studies: (1) use GPT-4o-mini to rewrite the solutions from the original training set; and (2) follow DART-Math (Tong et al., 2024) to use DeepSeekMath-7B-RL (Shao et al., 2024) to generate solutions for the fused problems.
 The results are shown in Table 6.
 We can see that the performance of the model fine-tuned on the rewritten solutions is better than the Standard setting, especially on the MATH and GSM8K datasets.
-However, the average improvement is only 1.3 points.
-Meanwhile, each fusion strategy of MathFusion still outperforms the rewritten solution by a large margin, indicating that the performance improvement of MathFusion mainly comes from the fusion of problems rather than the stronger teacher model.
-Appendix E Significant Test
+However, the average improvement is only 2.9 points.
+Meanwhile, each fusion strategy of MathFusion still outperforms the rewritten solution by a large margin.
+Additionally, though DeepSeekMath-7B-RL underperforms GPT-4o-mini in distillation quality (34.6 v.s.
+36.1 on average), it still outperforms DART-Math (34.6 v.s.
+32.9 on average).
+These results indicate that the performance improvement of MathFusion mainly comes from the new problem generated by three fusion strategies rather than the stronger teacher model.
+Appendix E Additional Baseline
+A most recent work, Mosaic-IT (Li et al., 2024c), shares similar idea with our MathFusion.
+Mosaic-IT is a model-free data augmentation technique that operates by concatenating existing instruction-following datasets and subsequently training LLMs using these augmented data instances along with meta-instructions.
+We conduct a comparison with the “Primary Strategy” proposed in Mosaic-IT, where the question pairs (same as MathFusion) and corresponding solutions from the original GSM8K and MATH datasets are concatenated into a single sample for SFT, resulting in 15K data.
+To mitigate overfitting to the pattern of answering multiple questions jointly, we also conduct an additional experiment which combine Mosaic-IT and the original GSM8K and MATH training sets during training, resulting in 30K data in total.
+The results are shown in Table 6.
+We observe that the Mosaic-IT leads to inferior performance, even worse than the Standard setting (i.e., using only the original GSM8K and MATH training data).
+We suspect this may be due to the lack of logical integration between problems when simply concatenated—unlike MathFusion, which explicitly introduces semantic or reasoning connections (e.g., sequential dependency or comparative logic) through its fusion strategies.
+This highlights the advantage of model-driven, structure-aware fusion over model-free concatenation.
+Appendix F Significant Test
 We conduct error analysis on MathFusion on Llama3-8B model to verify the consistent performance improvement of our MathFusionQA.
@@
 We also conduct statistical significance tests using the paired t-test, and results show that the performance improvement of MathFusion is statistically significant () on all benchmarks.
-Appendix F More Cases
+Appendix G More Cases
 More cases, including the original problems and , the fused problem , are shown below.
-Specifically, we show three reasonable cases in Case F, Case F, and Case F, and three unreasonable cases in Case F, Case F, and Case F.
+Specifically, we show three reasonable cases in Case G, Case G, and Case G, and three unreasonable cases in Case G, Case G, and Case G.